# OpenPBR: Novel Features and Implementation Details

### v1

@@ Abstract
The OpenPBR project originated in 2023 as a collaboration between Autodesk and Adobe to consolidate and refine their existing shading models, the Autodesk Standard Surface [Georgiev2019] and Adobe Standard Material [Kutz2021], respectively. The goal was to create a single, open, and standardized physically based shading model that would be suitable for the most common use cases in media and entertainment applications, while also being practical for real-time rendering.
OpenPBR is not intended to be a general framework for building material models (which exists in various forms [Gritz2010, Kettner2015, Smythe2016, Hillaire2023]), but a reasonably unambiguous description of a particular form of pre-built “uber-shader”, which experience has shown is practically very useful.
There is a long history of previous proposals for such a physically based uber-shader. Figure 2 shows an approximate genealogy, including Disney’s “Principled” shader [Burley2012], Allegorithmic’s PBR shading model [McDermott2018], and most recently Autodesk Standard Surface [Georgiev2019] and Adobe Standard Material [Kutz2021]. These most recent models have been found to be generally useful and practical for media and entertainment, so we based OpenPBR on them.
@@
Another positive outcome of defining a practical general-purpose model is the assurance that the existing standard general material frameworks [Gritz2010, Kettner2015, Smythe2016, Hillaire2023] have the required capabilities to implement it, and if not, arguing for and ensuring their adoption. In several cases, the effort has stimulated research into refinements of existing models, and generally helped to identify areas where further research on the underlying BSDF models is warranted.
@@ 2.1 Model formalism
While writing down this formalism is arguably not strictly necessary, it does help to clarify the structure of the model and how it is intended to be implemented in general material frameworks, and it also provides a common language for implementers to discuss the model. It was helpful to resolve potential ambiguities in the intended physical structure and meaning of the parameters, and to ensure that the model is well-defined in terms of light transport through the structure. Our model was designed explicitly to be compatible and consistent with the MaterialX framework [Smythe2016] (indeed, we contributed various changes to MaterialX to ensure compatibility), so the formalism we describe is therefore rather similar to what is provided in MaterialX.
@@
A slab does not itself specify anything about other slabs in relation to itself (e.g., its substrate slab or overlying slab). The adjacent medium above and below the slab will depend on where it sits in the eventual layer structure. The ambient dielectric medium at the very top of the entire structure (and bottom if thin-walled) is also assumed to be given and unspecified by the model. If the renderer keeps track of the dielectric medium in which the surface is embedded (via a scheme such as “nested dielectrics” [Schmidt:2002:SND]) – which may be the interior dielectric bulk of some transparent object in the scene, such as a piece of glass or body of water – then the surrounding ambient medium is a dielectric whose IOR we denote $n_{\mathrm{ambient}}$. Alternatively, if dielectric medium tracking is not performed, then $n_{\mathrm{ambient}}$ can be assumed to be 1 corresponding to air or vacuum.
@@
More generally, for transmissive BSDFs there is a similar constraint $E(\omega_{o})+T(\omega_{o})\leq 1$, where the directional transmittance $T(\omega_{o})$ is defined as the integral of the BSDF over the opposite hemisphere $\mathcal{H}_{-}$ to $\omega_{o}$. All BSDFs used are energy conserving in this sense. Additionally, a BSDF that does not dissipate any energy, such as a dielectric interface or a diffuse surface with a white albedo, is said to be energy preserving and satisfies $E(\omega_{o})+T(\omega_{o})=1$. (A technicality to be noted is that in order for the stated energy preservation to hold, the light transport has to be done in terms of the so-called “basic radiance“ $L/n^{2}$, where $n$ is the local index of refraction. This accounts for the squeezing of light rays as they enter a medium with a higher index of refraction. This is accounted for in standard derivations of BSDFs and implementation of light transport [Walter2007, Pharr2023].)
Reciprocity is the requirement that the BSDF is symmetric under interchange of the arguments, i.e., $\omega_{i}\leftrightarrow\omega_{o}$. [^2] Technically, the reciprocity condition also involves the index of refraction [Veach1998]. In some cases this is relaxed and a non-reciprocal BSDF or formula is used for simplicity or efficiency, as in practice it does not lead to serious issues (particularly if only doing unidirectional path tracing).
@@
In OpenPBR, we are mostly concerned with defining the layer structure rather than requiring any particular method be used to solve for the light transport, the exact solution of which is supposed to correspond to the desired ground truth. However, for practical purposes, it is important to give some suggested approximate schemes, at least as a starting point. In practice, even rather crude approximations can be acceptable if they yield plausible results. There is a large body of work in the graphics literature on modeling and rendering of layered material structures [Hanrahan1993, Ashikhmin2000, Kelemen2001, Weidlich2007, Jakob2014, Guo2018, Zeltner2018, Belcour2018]. Accurately solving for the light transport through such layered structures is a complex problem, which in general requires solving the radiative transfer equation (RTE) accounting for the scattering at the interfaces as well as inside the internal volumetric media. This is a non-trivial task that in general requires numerical methods such as Monte Carlo integration, or solving the RTE via various approximations such as diffusion theory or discrete ordinates methods [Pharr2023].
@@
In this spirit, the most simple representation of a layered configuration amounts to simply making a linear combination of the interface BSDFs $f_{\mathrm{sub}}$ and $f_{\mathrm{coat}}$. A common approach is the so-called albedo-scaling approximation [Smythe2016], where the total BSDF of this layered configuration is given by summing $f_{\mathrm{coat}}$ and $f_{\mathrm{sub}}$, with the substrate lobe weighted by a factor depending on the directional reflectance of the coat, which is designed to ensure that the resulting BSDF is energy conserving:
@@
It also ensures that if the substrate BSDF perfectly preserves energy (i.e., $E_{\mathrm{sub}}(\omega_{o})=1$) then the layer BSDF does also, ensuring that a “white furnace” test would pass. It ignores the physical requirement of reciprocity but captures the essential view dependence of the coat and substrate BSDFs, and is a good approximation for many practical cases. The Autodesk Standard Surface model [Georgiev2019], for example, uses this formulation for its layering.
@@
which, for example, is the formula used in the Autodesk Standard Surface [Georgiev2019] coat layer.
@@ 2.2 Microfacet models
The BSDFs $f_{\mathrm{conductor}}$, $f_{\mathrm{dielectric}}$, $f_{\mathrm{coat}}$ and $f_{\mathrm{diffuse}}$ of the metal, dielectric, coat and glossy-diffuse slabs, respectively, are each assumed to be described by a standard microfacet model. This is a widely used physical picture [Heitz2014, Pharr2023] in which the surface is assumed to be composed of a heightfield consisting of an ensemble of smooth microfacets of either metal, dielectric or Lambertian material, where the statistical distribution of the normal of these facets, termed the micronormal, determines the surface roughness characteristics at the macroscopic scale. (The fuzz model is distinct and based on a volumetric “microflake” model [Heitz2015].) At the present time, microfacet models are effectively the lowest-level physical description in the CG literature of the surface properties of materials. (There is interesting work on more sophisticated proposed formalisms that, for example, account for wave optical effects [Steinberg2022, Steinberg2024], but these are not yet widely used in practice.)
A microfacet BRDF has the standard single-scattering form [Walter2007, Pharr2023] consisting of a product of the Fresnel factor, the masking-shadowing function, and the Normal Distribution Function (NDF):  [^4] Also a Jacobian factor not shown here.
@@
where $h=(\omega_{i}+\omega_{o})/|\omega_{i}+\omega_{o}|$ is the half-vector (i.e., the micronormal), which mirror reflects $\omega_{i}$ into $\omega_{o}$. For dielectrics there is also a BTDF, i.e., the portion of the BSDF where the input and output directions lie in opposite rather than the same hemispheres, which has a similar form to the BRDF except with a modified half-vector, Fresnel factor, and Jacobian [Walter2007].
The Fresnel factor $F(\omega_{i},h)$ is determined by the complex index of refraction (IOR) of the reflecting material of each microfacet (or, more technically, the ratio of this to the exterior IOR), and its form differs depending on whether the material is a dielectric or conductor [Walter2007].
The masking-shadowing function $G(\omega_{i},\omega_{o})$ accounts for the probability that the input and output directions are occluded by the microsurface. It is usually derived using the Smith model, which determines $G$ given the NDF, and for the GGX NDF Equation 17 the masking-shadowing function then has a well-known form [Heitz2014].
The Normal Distribution Function (NDF) $D(m)$ describes the relative probability of occurrence of micronormal $m$ on the surface, and thus the roughness characteristics. We assume that the NDF is the so-called GGX distribution, which closely approximates the roughness of real materials (the name GGX derives from “ground glass”, but the formula was originally due to Trowbridge and Reitz [Trowbridge1975, Walter2007, Burley2012, Heitz2014, Pharr2023]), which has the basic form
@@
In practice, we restrict to the range $\alpha\in[0,1]$, as $\alpha>1$ does not produce a plausible rough appearance. Following Disney’s “Principled” shader [Burley2012], we set (in the isotropic case)
@@
The single-scattering microfacet BRDF of Equation 16 does not conserve energy, as it neglects to account for multiple scattering between the microfacets. In OpenPBR, we generally consider it important that a material which is supposed to be non-absorbing (e.g., a metal with Fresnel factor close to 1) should conserve energy and pass a white furnace test. Thus an implementation should account for multiple scattering via one of a number of schemes, otherwise the reflection from rough metals and dielectrics will be dimmer and less saturated than it should be. An accurate approach based on the Smith microsurface model is described in [Heitz2016a]. Simpler approximate models are presented in [Kulla2017] (which functions by adding compensation lobes to account for the missing energy), and [Turquin2019] (which scales the albedo of the lobe to maintain energy preservation at the expense of reciprocity).
@@
which reduces to the isotropic form when $\alpha_{t}=\alpha_{b}=\alpha$. Efficient techniques for sampling BSDFs employing the anisotropic GGX microfacet model are presented in [Heitz2018, Dupuy2023].
@@ 2.3 Model structure
For convenience and efficiency, at present it is most likely to be mapped to a model consisting of a mixture of BSDF lobes similar to the Autodesk Standard Surface shader [Georgiev2019] and its representation in MaterialX. An example derivation of such a model is provided in Appendix B. In the official repository we provide a reference implementation in MaterialX, which has a structure similar to this derivation.
@@ 3.1.1 Glossy-diffuse model: EON
There is a long history in computer graphics of modeling such glossy-diffuse, plastic-like materials via a combination of a specular term and diffuse term, with heuristics to ensure energy conservation (Figure 16). The most well-known example is the Ashikhmin–Shirley model [Ashikhmin2000, Pharr2023], which uses a microfacet specular term and a diffuse term based on the Lambertian BRDF.
@@
where $f_{\mathrm{diffuse}}$ is based on the classic Oren–Nayar model [OrenNayar1994], which is a V-cavity based microfacet model in which the microfacets are assumed to be individually Lambertian. The Oren–Nayar model is parameterized by roughness $\sigma$ and an overall RGB scale factor $\boldsymbol{\rho}$. The roughness models a “flattening” effect (Figure 17), which is observed in real rough diffuse materials.
During our investigation, we found that the original Oren–Nayar model loses a significant amount of energy when the roughness is high, which is not compatible with the energy-preservation requirements of the OpenPBR model. To address this, we developed a modified version of the Oren–Nayar model (termed the “energy-preserving Oren–Nayar model”, or EON model) that preserves energy at high roughness values, as described by [Portsmouth2025]. This augments the Oren–Nayar model with a simple analytical, reciprocal energy compensation term:
@@
The Oren–Nayar term $f_{\mathrm{ON}}$ is given by the form (introduced by Fujii [Portsmouth2025])
@@
where the albedo of the Oren–Nayar lobe $f_{\mathrm{diffuse}}(\omega_{i},\omega_{o})$ must be taken to be $\mathbf{C}=\mathtt{base\_weight}\times\mathtt{base\_color}$ in order to satisfy Equation 40. If a reciprocal formulation is desired, the classic Ashikhmin–Shirley or Kelemen model [Ashikhmin2000, Kelemen2001, Kulla2017, Kutz2021] form is
@@
where $\mathcal{N}$ is a normalization factor such that Equation 40 is satisfied. In the case of a zero roughness (i.e., Lambertian) Oren–Nayar base $\mathcal{N}$ can be tabulated in terms of the dielectric IOR and roughness [Kutz2021], and the required albedo of the Lambertian base is equal to $\mathbf{C}$ as in the non-reciprocal albedo-scaling approximation. Extending this to the more general case of a non-Lambertian rough Oren–Nayar base would require adding the roughness dimension to the tabulation, and the required Oren–Nayar albedo will not simply equal $\mathbf{C}$. We leave the specific choice of model and these details to the implementation.
The EON diffuse model also provides more sophisticated importance sampling than standard cosine-weighted sampling typically used for diffuse materials [Portsmouth2025]. The sampling is based on Linearly Transformed Cosine (LTC) sampling [Heitz2016b], with a modification to avoid generating redundant samples under the hemisphere (so-called CLTC sampling). This results in a significant reduction in noise (see Figures 18 and  19).

### v2

@@ Abstract
The OpenPBR project originated in 2023 as a collaboration between Autodesk and Adobe to consolidate and refine their existing shading models, the Autodesk Standard Surface [17] and Adobe Standard Material [37], respectively. The goal was to create a single, open, and standardized physically based shading model that would be suitable for the most common use cases in media and entertainment applications, while also being practical for real-time rendering.
OpenPBR is not intended to be a general framework for building material models (which exists in various forms [18, 34, 46, 29]), but a reasonably unambiguous description of a particular form of pre-built “uber-shader”, which experience has shown is practically very useful.
There is a long history of previous proposals for such a physically based uber-shader. Figure 2 shows an approximate genealogy, including Disney’s “Principled” shader [10], Allegorithmic’s PBR shading model [38], and most recently Autodesk Standard Surface [17] and Adobe Standard Material [37]. These most recent models have been found to be generally useful and practical for media and entertainment, so we based OpenPBR on them.
@@
Another positive outcome of defining a practical general-purpose model is the assurance that the existing standard general material frameworks [18, 34, 46, 29] have the required capabilities to implement it, and if not, arguing for and ensuring their adoption. In several cases, the effort has stimulated research into refinements of existing models, and generally helped to identify areas where further research on the underlying BSDF models is warranted.
@@ 2.1 Model formalism
While writing down this formalism is arguably not strictly necessary, it does help to clarify the structure of the model and how it is intended to be implemented in general material frameworks, and it also provides a common language for implementers to discuss the model. It was helpful to resolve potential ambiguities in the intended physical structure and meaning of the parameters, and to ensure that the model is well-defined in terms of light transport through the structure. Our model was designed explicitly to be compatible and consistent with the MaterialX framework [46] (indeed, we contributed various changes to MaterialX to ensure compatibility), so the formalism we describe is therefore rather similar to what is provided in MaterialX.
@@
A slab does not itself specify anything about other slabs in relation to itself (e.g., its substrate slab or overlying slab). The adjacent medium above and below the slab will depend on where it sits in the eventual layer structure. The ambient dielectric medium at the very top of the entire structure (and bottom if thin-walled) is also assumed to be given and unspecified by the model. If the renderer keeps track of the dielectric medium in which the surface is embedded (via a scheme such as “nested dielectrics” [45]) – which may be the interior dielectric bulk of some transparent object in the scene, such as a piece of glass or body of water – then the surrounding ambient medium is a dielectric whose IOR we denote $n_{\mathrm{ambient}}$. Alternatively, if dielectric medium tracking is not performed, then $n_{\mathrm{ambient}}$ can be assumed to be 1 corresponding to air or vacuum.
@@
More generally, for transmissive BSDFs there is a similar constraint $E(\omega_{o})+T(\omega_{o})\leq 1$, where the directional transmittance $T(\omega_{o})$ is defined as the integral of the BSDF over the opposite hemisphere $\mathcal{H}_{-}$ to $\omega_{o}$. All BSDFs used are energy conserving in this sense. Additionally, a BSDF that does not dissipate any energy, such as a dielectric interface or a diffuse surface with a white albedo, is said to be energy preserving and satisfies $E(\omega_{o})+T(\omega_{o})=1$. (A technicality to be noted is that in order for the stated energy preservation to hold, the light transport has to be done in terms of the so-called “basic radiance“ $L/n^{2}$, where $n$ is the local index of refraction. This accounts for the squeezing of light rays as they enter a medium with a higher index of refraction. This is accounted for in standard derivations of BSDFs and implementation of light transport [52, 41].)
Reciprocity is the requirement that the BSDF is symmetric under interchange of the arguments, i.e., $\omega_{i}\leftrightarrow\omega_{o}$. [^2] Technically, the reciprocity condition also involves the index of refraction [51]. In some cases this is relaxed and a non-reciprocal BSDF or formula is used for simplicity or efficiency, as in practice it does not lead to serious issues (particularly if only doing unidirectional path tracing).
@@
In OpenPBR, we are mostly concerned with defining the layer structure rather than requiring any particular method be used to solve for the light transport, the exact solution of which is supposed to correspond to the desired ground truth. However, for practical purposes, it is important to give some suggested approximate schemes, at least as a starting point. In practice, even rather crude approximations can be acceptable if they yield plausible results. There is a large body of work in the graphics literature on modeling and rendering of layered material structures [23, 2, 33, 53, 31, 22, 57, 7]. Accurately solving for the light transport through such layered structures is a complex problem, which in general requires solving the radiative transfer equation (RTE) accounting for the scattering at the interfaces as well as inside the internal volumetric media. This is a non-trivial task that in general requires numerical methods such as Monte Carlo integration, or solving the RTE via various approximations such as diffusion theory or discrete ordinates methods [41].
@@
In this spirit, the most simple representation of a layered configuration amounts to simply making a linear combination of the interface BSDFs $f_{\mathrm{sub}}$ and $f_{\mathrm{coat}}$. A common approach is the so-called albedo-scaling approximation [46], where the total BSDF of this layered configuration is given by summing $f_{\mathrm{coat}}$ and $f_{\mathrm{sub}}$, with the substrate lobe weighted by a factor depending on the directional reflectance of the coat, which is designed to ensure that the resulting BSDF is energy conserving:
@@
It also ensures that if the substrate BSDF perfectly preserves energy (i.e., $E_{\mathrm{sub}}(\omega_{o})=1$) then the layer BSDF does also, ensuring that a “white furnace” test would pass. It ignores the physical requirement of reciprocity but captures the essential view dependence of the coat and substrate BSDFs, and is a good approximation for many practical cases. The Autodesk Standard Surface model [17], for example, uses this formulation for its layering.
@@
which, for example, is the formula used in the Autodesk Standard Surface [17] coat layer.
@@ 2.2 Microfacet models
The BSDFs $f_{\mathrm{conductor}}$, $f_{\mathrm{dielectric}}$, $f_{\mathrm{coat}}$ and $f_{\mathrm{diffuse}}$ of the metal, dielectric, coat and glossy-diffuse slabs, respectively, are each assumed to be described by a standard microfacet model. This is a widely used physical picture [27, 41] in which the surface is assumed to be composed of a heightfield consisting of an ensemble of smooth microfacets of either metal, dielectric or Lambertian material, where the statistical distribution of the normal of these facets, termed the micronormal, determines the surface roughness characteristics at the macroscopic scale. (The fuzz model is distinct and based on a volumetric “microflake” model [24].) At the present time, microfacet models are effectively the lowest-level physical description in the CG literature of the surface properties of materials. (There is interesting work on more sophisticated proposed formalisms that, for example, account for wave optical effects [48, 47], but these are not yet widely used in practice.)
A microfacet BRDF has the standard single-scattering form [52, 41] consisting of a product of the Fresnel factor, the masking-shadowing function, and the Normal Distribution Function (NDF):  [^4] Also a Jacobian factor not shown here.
@@
where $h=(\omega_{i}+\omega_{o})/|\omega_{i}+\omega_{o}|$ is the half-vector (i.e., the micronormal), which mirror reflects $\omega_{i}$ into $\omega_{o}$. For dielectrics there is also a BTDF, i.e., the portion of the BSDF where the input and output directions lie in opposite rather than the same hemispheres, which has a similar form to the BRDF except with a modified half-vector, Fresnel factor, and Jacobian [52].
The Fresnel factor $F(\omega_{i},h)$ is determined by the complex index of refraction (IOR) of the reflecting material of each microfacet (or, more technically, the ratio of this to the exterior IOR), and its form differs depending on whether the material is a dielectric or conductor [52].
The masking-shadowing function $G(\omega_{i},\omega_{o})$ accounts for the probability that the input and output directions are occluded by the microsurface. It is usually derived using the Smith model, which determines $G$ given the NDF, and for the GGX NDF Equation 17 the masking-shadowing function then has a well-known form [27].
The Normal Distribution Function (NDF) $D(m)$ describes the relative probability of occurrence of micronormal $m$ on the surface, and thus the roughness characteristics. We assume that the NDF is the so-called GGX distribution, which closely approximates the roughness of real materials (the name GGX derives from “ground glass”, but the formula was originally due to Trowbridge and Reitz [49, 52, 10, 27, 41]), which has the basic form
@@
In practice, we restrict to the range $\alpha\in[0,1]$, as $\alpha>1$ does not produce a plausible rough appearance. Following Disney’s “Principled” shader [10], we set (in the isotropic case)
@@
The single-scattering microfacet BRDF of Equation 16 does not conserve energy, as it neglects to account for multiple scattering between the microfacets. In OpenPBR, we generally consider it important that a material which is supposed to be non-absorbing (e.g., a metal with Fresnel factor close to 1) should conserve energy and pass a white furnace test. Thus an implementation should account for multiple scattering via one of a number of schemes, otherwise the reflection from rough metals and dielectrics will be dimmer and less saturated than it should be. An accurate approach based on the Smith microsurface model is described in [26]. Simpler approximate models are presented in [35] (which functions by adding compensation lobes to account for the missing energy), and [50] (which scales the albedo of the lobe to maintain energy preservation at the expense of reciprocity).
@@
which reduces to the isotropic form when $\alpha_{t}=\alpha_{b}=\alpha$. Efficient techniques for sampling BSDFs employing the anisotropic GGX microfacet model are presented in [28, 14].
@@ 2.3 Model structure
For convenience and efficiency, at present it is most likely to be mapped to a model consisting of a mixture of BSDF lobes similar to the Autodesk Standard Surface shader [17] and its representation in MaterialX. An example derivation of such a model is provided in Appendix B. In the official repository we provide a reference implementation in MaterialX, which has a structure similar to this derivation.
@@ 3.1.1 Glossy-diffuse model: EON
There is a long history in computer graphics of modeling such glossy-diffuse, plastic-like materials via a combination of a specular term and diffuse term, with heuristics to ensure energy conservation (Figure 16). The most well-known example is the Ashikhmin–Shirley model [2, 41], which uses a microfacet specular term and a diffuse term based on the Lambertian BRDF.
@@
where $f_{\mathrm{diffuse}}$ is based on the classic Oren–Nayar model [39], which is a V-cavity based microfacet model in which the microfacets are assumed to be individually Lambertian. The Oren–Nayar model is parameterized by roughness $\sigma$ and an overall RGB scale factor $\boldsymbol{\rho}$. The roughness models a “flattening” effect (Figure 17), which is observed in real rough diffuse materials.
During our investigation, we found that the original Oren–Nayar model loses a significant amount of energy when the roughness is high, which is not compatible with the energy-preservation requirements of the OpenPBR model. To address this, we developed a modified version of the Oren–Nayar model (termed the “energy-preserving Oren–Nayar model”, or EON model) that preserves energy at high roughness values, as described by [43]. This augments the Oren–Nayar model with a simple analytical, reciprocal energy compensation term:
@@
The Oren–Nayar term $f_{\mathrm{ON}}$ is given by the form (introduced by Fujii [43])
@@
where the albedo of the Oren–Nayar lobe $f_{\mathrm{diffuse}}(\omega_{i},\omega_{o})$ must be taken to be $\mathbf{C}=\mathtt{base\_weight}\times\mathtt{base\_color}$ in order to satisfy Equation 40. If a reciprocal formulation is desired, the classic Ashikhmin–Shirley or Kelemen model [2, 33, 35, 37] form is
@@
where $\mathcal{N}$ is a normalization factor such that Equation 40 is satisfied. In the case of a zero roughness (i.e., Lambertian) Oren–Nayar base $\mathcal{N}$ can be tabulated in terms of the dielectric IOR and roughness [37], and the required albedo of the Lambertian base is equal to $\mathbf{C}$ as in the non-reciprocal albedo-scaling approximation. Extending this to the more general case of a non-Lambertian rough Oren–Nayar base would require adding the roughness dimension to the tabulation, and the required Oren–Nayar albedo will not simply equal $\mathbf{C}$. We leave the specific choice of model and these details to the implementation.
The EON diffuse model also provides more sophisticated importance sampling than standard cosine-weighted sampling typically used for diffuse materials [43]. The sampling is based on Linearly Transformed Cosine (LTC) sampling [25], with a modification to avoid generating redundant samples under the hemisphere (so-called CLTC sampling). This results in a significant reduction in noise (see Figures 18 and  19).